OpenStudy (anonymous):

Solve |2x - 5| = 4. {x | x = -4.5 or x = 4.5} {x | x = 0.5 or x = 4.5} {x | 0.5 < x < 4.5}

OpenStudy (bagelini):

are these all different problems? or are you telling us what we replace "x" with?

OpenStudy (anonymous):

|4| does thtat equal four or -4

OpenStudy (bagelini):

When something has the absolute value brackets: \[\left| \right|\] it is talking about how far away the number is from zero on a number line.

OpenStudy (anonymous):

so -4

OpenStudy (bagelini):

so in your case, \[\left| -4 \right|\] would still be +4 because it is four points away from zero on a number line. does that make sense?

OpenStudy (anonymous):

i plugged a and b in and they both dont work so i think its c can you check??

OpenStudy (anonymous):

its b or c actually im not shure

OpenStudy (mathstudent55):

To solve an absolute value equation like yours, solve the compound equation below: 2x - 5 = 4 or 2x - 5 = -4

OpenStudy (mathstudent55):

Solve each simple equation and keep the word "or" between solutions. That is you final answer.

OpenStudy (anonymous):

isnt ther no soulution to thta tho

OpenStudy (mathstudent55):

Can you solve 2x - 5 = 4?

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

so its a ????

OpenStudy (mathstudent55):

You are correct. x = 4.5 for the left equation. Now do the right equation. 2x - 5 = 4 or 2x - 5 = -4 Add 5 to both sides of both equations: 2x = 9 or 2x = 1 Divide both sides of both equations by 2. x = 4.5 or x = 0.5

OpenStudy (anonymous):

so its c

OpenStudy (mathstudent55):

No. This problem has no greater than or less than. It is strictly an equation. The answer is b.